OpenStudy (anonymous):

What is the period and frequency of a 10-pound weight oscillating (as a pendulum) on a 5-meter-long rope?

OpenStudy (anonymous):

frequency=2pi/omega=2pi/sqrt(g/l)=2pi/sqrt(10/5)=2pi/sqrt(2) Hz

OpenStudy (anonymous):

For a pendulum: T(period)=\[2\pi \sqrt{l/g}\] where l is the length of the rope and g is gravity. f(frequency)=\[1/T\] Therefore assuming that this is on earth the period is 4.487989505 and the frequency is 0.228169203.